# Correcting for sampling variability in maximum likelihood-based one-sample log-rank tests

### v1

@@ Abstract
Single-arm studies in the early development phases of new treatments are not uncommon in the context of rare diseases or in paediatrics. If an assessment of efficacy is to be made at the end of such a study, the observed endpoints can be compared with reference values that can be derived from historical data. For a time-to-event endpoint, a statistical comparison with a reference curve can be made using the one-sample log-rank test. In order to ensure the interpretability of the results of this test, the role of the reference curve is crucial. This quantity is often estimated from a historical control group using a parametric procedure. Hence, it should be noted that it is subject to estimation uncertainty. However, this aspect is not taken into account in the one-sample log-rank test statistic. We analyse this estimation uncertainty for the common situation that the reference curve is estimated parametrically using the maximum likelihood method, and indicate how the variance estimation of the one-sample log-rank test can be adapted in order to take this variability into account. The resulting test procedures are illustrated using a data example and analysed in more detail using simulations.
@@ 2 Notation
We are interested in the distribution of a time-to-event variable $T$. We assume that the observations in both groups are distributed as the two random variables $T_{A}$ and $T_{B}$, respectively. However, observation of this variable might be censored due to an independent censoring mechanism through the random variable $C_{A}$ in the historic control group or $C_{B}$ in the prospective group. In any case, we can only observe $X_{x}\coloneqq\min(T_{x},C_{x})$ and $\delta_{x}\coloneqq\mathbbm{1}(T_{x}\leq C_{x})$ for $x\in\{A,B\}$. We assume that all patients are observed for a maximum time $t_{\text{max}}<\infty$, i.e. $X\leq t_{\text{max}}$ with probability 1. Density, cumulative distribution, survival, hazard and cumulative hazard functions will be denoted by $f,F,S,\lambda$ and $\Lambda$, respectively. The variable to which the function refers is specified in the index. For example, $f_{T_{A}}$ denotes the density function of the random variable $T_{A}$, i.e. the time-to-event variable in the historical control cohort. For $x\in\{A,B\}$, we assume that $n_{x}$ independent and identically distributed replicates of the tuple $(X_{x},\delta_{x})$ are given which we will denote by $(X_{x,i},\delta_{x,i})$ for $i\in\{1,\dots,n_{x}\}$. We also assume that the data from both groups is independent of each other. By the functions $N_{x,i},Y_{x,i}\colon\mathbb{R}_{+}\to\{0,1\}$ we denote the time-dependent event and at-risk indicators for patient $i\in\{1,\dots,n_{x}\}$ in group $x\in\{A,B\}$, i.e.
@@ 6 Simulation study
The aim of this simulation study is to quantify the type I error rate inflation of the one-sample log-rank test when it is interpreted as a test of the null hypothesis $H_{0}$ instead of $H_{0,\text{OSLR}}$ in settings where the reference curve is given by a parametric estimate from historic control data. At the same time, we want to investigate whether the corrected test, that accounts for the additional variability $V_{2}$ introduced by the estimation procedure, maintains the nominal type I error rate.
For the historic control arm as well as for the experimental arm, we assume that the patients were recruited uniformly between the start and the end of the accrual period after $a=2$ years. Each patient remains in the trial until the event of interest occurs or the end of the follow-up period of length $f=3$ years that follows the accrual period. Hence, we have $C_{x}\sim\mathcal{U}[f,a+f]$ for $x\in\{A,B\}$. The recruitment date and hence also the date of administrative censoring are simulated independently from the time of the event of interest. As our aim is to analyse the adherence to the nominal type I error rate, we use the same distribution for $T_{A}$ and $T_{B}$. We assume a Weibull distribution with shape parameter $\kappa\in\{0.5,1,2\}$ with a fixed one-year survival rate of 50%. Hence, the cumulative hazard function at time $s$ is given by $\Lambda_{T_{A}}(s)=\Lambda_{T_{B}}(s)=-\log(0.5)\cdot s^{\kappa}$.
@@
For each combination of those parameters we conduct 100,000 simulation runs. In each run, we use the R package flexsurv to fit a Weibull distribution to the data from group $A$ and conduct the standard one-sample log-rank test to compare the data from group B with the resulting reference curve. At the same time, we also compute the corrected test statistics and the two-sample log-rank test statistic comparing the two groups directly. The Monte Carlo error interval for 100,000 simulation runs for the nominal (one-sided) rate of 0.025 is $[0.0240;0.0260]$. For the (two-sided) rate of 0.05 it amounts to $[0.0486;0.0514]$.
@@
Uncorrected OSLR One-sample log-rank test with parametrically estimated reference curve, but without correction for additional variability.
Uncorrected OSLR (Wu) As above, but with variance estimation according to [19], i.e. $w=0.5$ in (5).
Corrected OSLR One-sample log-rank test with parametrically estimated reference curve and corrected variance estimation.
Corrected OSLR (Wu) As above, but with variance estimation for the OSLR contribution according to [19], i.e. $w=0.5$ in (5).
@@
In Figure 2 the two-sided empirical type I error rates for all procedures are shown for $\kappa=1$. On the left side, the sample size of the experimental group is fixed at $n_{B}=50$. Obviously, the two uncorrected procedures that use the parametrically estimated reference curve do not at all comply with the nominal type I error level when regarded as a test of $H_{0}$. As in [5], the discrepancy increases drastically as the allocation ratio increases. Within the considered range of allocation ratios, the increase appears to be almost linear in the allocation ratio. All other procedures match to a greater or lesser extent the nominal rate. This includes the two procedures with the additional correction for the application of the estimated reference curve. However, probably due to small sample sizes, not all empirical rates lie within the (grey shaded) confidence interval. This is also the case for the standard testing procedure for two-sample survival data, the two-sample log-rank test. On the right hand side of Figure 2, the allocation ratio is fixed at $\pi=0.5$ and the sample size is varied. The type I error rate inflation of the uncorrected procedures is almost not affected by this variation. This resembles the results in [5] and can also be explained by our theory as none of the factors of the additional variance $V_{2}$ in (6) is affected by the variation of the sample size while the allocation ratio is held constant. However, for the uncorrected as well as for the corrected procedures and the two-sample log-rank test we can observe a slightly larger inflation for small sample sizes. This may be because the asymptotics are not yet taking effect here.
@@
Similar observations can be made for the one-sided rates shown in Figure 3. These rates show the rejection rates of $H_{0}$ in favour of the experimental treatment. The choice of the weight $w$ in (5) plays a much more important role here. While the rejection rates for the procedures using $w=0$ and $w=0.5$ were almost indistinguishable in Figure 2, this is not the case here anymore. As also found in [19, 5] the standard variance estimation leads to strongly anti-conservative tests when trying to prove superiority of the new treatment. Choosing $w=0.5$ as in [19] yields better adherence to the nominal level. In particular, the corrected one-sample log-rank test performs quite well for unbalanced sample sizes but the performance deteriorates when the allocation ratio increases and the overall sample size decreases. This development is contrary to that for the standard two-sample test. As also shown in [11] it is conservative in unbalanced settings. However, performance improved if the sample size of the control cohort decreases.
Results for other settings are very similar to those shown here. Corresponding plots can be found in the Supplementary Material.
@@ 7 Discussion
As we have demonstrated here, a considerable inflation of the type I error rate occurs if the one-sample log-rank test is to be interpreted as a test of the practically highly relevant null hypothesis $H_{0}$. This is the case as the calculation of the test statistic does not take the estimation uncertainty of the reference curve into account. With the analytical derivation and a simulation study, we were able to demonstrate that we can decompose the variance of the test statistic under this hypothesis into two parts, one of which corresponds to the well-known variance estimators for the one-sample log-rank test and the other addresses the uncertainty in the reference curve. As can be seen from formula (7), the allocation ratio between the historical control cohort and the experimental cohort is of fundamental importance. While the one-sample log-rank test in its current form can still play an important role as a helpful tool, particularly in the early stages of treatment development, especially in rare diseases or in paediatrics, we would like to draw attention to the fact that there are ways of taking a more critical statistical view of the role of the reference curve. For example, the test statistics developed here could be considered as an additional sensitivity analysis to the one-sample log-rank test.
Of course, it can be argued that the resulting comparison is ultimately a two-group comparison between the historical cohort and the study cohort, which could also be achieved with a two-sample log-rank test. Nevertheless, our proposed method has the advantage that we could achieve efficiency gains by estimating from the historical data using a parametric rather than a fully non-parametric method. This preserves efficiency gains that are urgently needed in scenarios where the one-sample log-rank test is frequently used. Compared to a purely parametric comparison of both groups, the test method presented here at least preserves the basic idea of the one-sample log-rank test and does not assume that the distribution in group B belongs to the same parametric family as the one selected for group A. Also, our proposed method could be used as an additional sensitivity analysis if the one-sample log-rank test shall still be applied for the primary analysis.
@@
Going beyond the standard one-sample log-rank test, the techniques presented and evaluated here are in general also applicable to variations of the one-sample log-rank test. This includes weighted tests [4], and the resulting combination tests [17] as well as multi-stage procedures [13, 16].
@@ Appendix A Derivation of the asymptotic distribution
$-n_{B}^{1/2}\left(\mathbb{E}[\Lambda(\hat{\boldsymbol{\theta}}_{A},t\wedge X_{B})|\hat{\boldsymbol{\theta}}_{A}]-\mathbb{E}_{X_{B}}[\Lambda(\boldsymbol{\theta}_{A},t\wedge X_{B})]\right)$ (17)

### v2

@@ Abstract
Single-arm studies in the early development phases of new treatments are not uncommon in the context of rare diseases or in paediatrics. If an assessment of efficacy is to be made at the end of such a study, the observed endpoints can be compared with reference values that can be derived from historical data. For a time-to-event endpoint, a statistical comparison with a reference curve can be made using the one-sample log-rank test. In order to ensure the interpretability of the results of this test, the role of the reference curve is crucial. This quantity is often estimated from a historical control group using a parametric procedure. Hence, it should be noted that it is subject to estimation uncertainty. However, this aspect is not taken into account in the one-sample log-rank test statistic. We analyse this estimation uncertainty for the common situation that the reference curve is estimated parametrically using the maximum likelihood method, and indicate how the variance estimation of the one-sample log-rank test can be adapted in order to take this variability into account. The resulting test procedures are illustrated using a data example and analysed in more detail using simulations, particularly in comparison with established two-sample methods.
@@ 2 Notation
We are interested in the distribution of a time-to-event variable $T$. We assume that the observations in both groups are distributed as the two random variables $T_{A}$ and $T_{B}$, respectively. However, observation of this variable might be censored due to an independent censoring mechanism through the random variable $C_{A}$ in the historic control group or $C_{B}$ in the prospective group. In any case, we can only observe $X_{x}\coloneqq\min(T_{x},C_{x})$ and $\delta_{x}\coloneqq\mathbbm{1}(T_{x}\leq C_{x})$ for $x\in\{A,B\}$. We assume that all patients are observed for a maximum time $t_{\text{max}}<\infty$, i.e. $X\leq t_{\text{max}}$ with probability 1. Density, cumulative distribution, survival, hazard and cumulative hazard functions will be denoted by $f,F,S,\lambda$ and $\Lambda$, respectively. The variable to which the function refers is specified in the index. For example, $f_{T_{A}}$ denotes the density function of the random variable $T_{A}$, i.e. the time-to-event variable in the historical control cohort. For $x\in\{A,B\}$, we assume that $n_{x}$ independent and identically distributed replicates of the tuple $(X_{x},\delta_{x})$ are given which we will denote by $(X_{x,i},\delta_{x,i})$ for $i\in\{1,\dots,n_{x}\}$. We also assume that the data from both groups are independent of each other. By the functions $N_{x,i},Y_{x,i}\colon\mathbb{R}_{+}\to\{0,1\}$ we denote the time-dependent event and at-risk indicators for patient $i\in\{1,\dots,n_{x}\}$ in group $x\in\{A,B\}$, i.e.
@@ 6 Simulation study
The aim of this simulation study is to quantify the type I error rate inflation of the one-sample log-rank test when it is interpreted as a test of the null hypothesis $H_{0}$ instead of $H_{0,\text{OSLR}}$ in settings where the reference curve is given by a parametric estimate from historic control data. At the same time, we want to investigate whether the corrected test, that accounts for the additional variability $V_{2}$ introduced by the estimation procedure, maintains the nominal type I error rate. Furthermore, we want to examine whether the corrected test has power advantages over the two-sample log-rank test due to the parametric assumption for the historical control cohort.
6.1 Data generation
For the historic control arm as well as for the experimental arm, we assume that the patients were recruited uniformly between the start and the end of the accrual period after $a=2$ years. Each patient remains in the trial until the event of interest occurs or the end of the follow-up period of length $f=3$ years that follows the accrual period. Hence, we have $C_{x}\sim\mathcal{U}[f,a+f]$ for $x\in\{A,B\}$. The recruitment date and hence also the date of administrative censoring are simulated independently from the time of the event of interest. As our aim is to analyse the adherence to the nominal type I error rate, we use the same distribution for $T_{A}$ and $T_{B}$. We assume a Weibull distribution for the historical control group with shape parameter $\kappa\in\{0.5,1,2\}$ with a fixed one-year survival rate of 50%. Hence, the cumulative hazard function at time $s$ is given by $\Lambda_{T_{A}}(s)=-\log(0.5)\cdot s^{\kappa}$. The cumulative hazard function for the experimental group is given by a proportional hazards model via $\Lambda_{T_{B}}(s)=\omega\cdot\Lambda_{T_{A}}(s)$. In order to investigate type I error rates as well as power, we considered considered hazard ratios $\omega\in\{1,0.8\}$. Please note that a Weibull distribution with shape parameter $\kappa=1$ corresponds to an exponential distribution.
@@ 6.1 Data generation
For each combination of those parameters we conduct 100,000 simulation runs. In each run, we use the R package flexsurv to fit a distribution to the data from group $A$ and conduct the standard one-sample log-rank test to compare the data from group $B$ with the resulting reference curve. At the same time, we also compute the corrected test statistics and the two-sample log-rank test statistic comparing the two groups directly. For all scenarios, we use Weibull MLE estimates for the procedures that require parametric estimates. If $\kappa=1$, the time-to-event endpoint is exponentially distributed. In these scenarios, we will also consider estimates for an exponential distribution. In this way, we can also uncover potential advantages of a more sparse parametric modeling of the refrence survival curve. The Monte Carlo error interval for 100,000 simulation runs for the nominal (one-sided) rate of 0.025 is $[0.0240;0.0260]$. For the (two-sided) rate of 0.05 it amounts to $[0.0486;0.0514]$.
@@
Uncorrected OSLR One-sample log-rank test with parametrically estimated reference curve, but without correction for additional variability. If $\kappa=1$, we also consider a version with estimation from exponential distribution.
Uncorrected OSLR (Wu) As above, but with variance estimation according to [19], i.e. $w=0.5$ in (5). If $\kappa=1$, we also consider a version with estimation from exponential distribution.
Corrected OSLR One-sample log-rank test with parametrically estimated reference curve and corrected variance estimation. If $\kappa=1$, we also consider a version with estimation from exponential distribution.
Corrected OSLR (Wu) As above, but with variance estimation for the OSLR contribution according to [19], i.e. $w=0.5$ in (5). If $\kappa=1$, we also consider a version with estimation from exponential distribution.
@@ 6.2.1 Type I error rate
In Figure 2 the two-sided empirical type I error rates for all procedures are shown for $\kappa=1$. On the left side, the sample size of the experimental group is fixed at $n_{B}=100$. Obviously, the two uncorrected procedures that use the parametrically estimated reference curve do not at all comply with the nominal type I error level when regarded as a test of $H_{0}$. As in [5], the discrepancy increases drastically as the allocation ratio increases. Within the considered range of allocation ratios, the increase appears to be almost linear in the allocation ratio. All other procedures match to a greater or lesser extent the nominal rate. This includes the two procedures with the additional correction for the application of the estimated reference curve. However, probably due to small sample sizes, not all empirical rates lie within the (grey shaded) Monte Carlo error interval. This is also the case for the standard testing procedure for two-sample survival data, the two-sample log-rank test. On the right hand side of Figure 2, the allocation ratio is fixed at $\pi=0.25$ and the sample size is varied. The type I error rate inflation of the uncorrected procedures is hardly affected by this variation. This resembles the results in [5] and can also be explained by our theory as none of the factors of the additional variance $V_{2}$ in (6) is affected by the variation of the sample size while the allocation ratio is held constant. However, for the uncorrected as well as for the corrected procedures and the two-sample log-rank test we can observe a slightly larger inflation for small sample sizes. This may be because the asymptotics are not yet taking effect here. As can be seen from the results in the Supplementary Material, there is no remarkable difference in these rates depending on whether an exponential or a Weibull distribution is assumed in the case of $\kappa=1$.
@@
Similar observations can be made for the one-sided rates shown in Figure 3. These rates show the rejection rates of $H_{0}$ in favour of the experimental treatment. The choice of the weight $w$ in (5) plays a much more important role here. While the rejection rates for the procedures using $w=0$ and $w=0.5$ were almost indistinguishable in Figure 2, this is not the case here anymore. As also found in [19, 5] the standard variance estimation leads to strongly conservative tests when trying to prove superiority of the new treatment. Choosing $w=0.5$ as in [19] yields a noticeably less conservative testing procedure. In particular, the corrected one-sample log-rank test performs quite well for unbalanced sample sizes but the performance deteriorates when the allocation ratio increases and the overall sample size decreases. This development is contrary to that for the standard two-sample test. As already shown in [11], the two-sample log-rank test is conservative in unbalanced settings. However, performance improved if the sample size of the control cohort decreases.
As with the two-sided rates, it makes no noticeable difference whether the more restrictive assumption of an exponential distribution is made for the case $\kappa=1$. Results for other settings are very similar to those shown here. Corresponding plots can be found in the Supplementary Material.
6.2.2 Power
As the results for $\omega=1$, i.e. under the null hypothesis $H_{0}$, have shown, the corrected versions of the one-sample log-rank test are indeed suitable as a test for this hypothesis. Consequently, we now want to evaluate whether the corrected test achieves efficiency gains by making a parametric assumption about one of the two groups instead of using a completely non-parametric method.
Figure 4: Empirical power for five procedures that showed acceptable adherence to the nominal type I error rate with fixed sample size in the experimental group and varying allocation ratios (left) or varying sample size in the experimental group with fixed allocation ratio (right).
For the hazard ratio $\omega=0.8$ between the two groups, Figure 4 shows the power of the corrected procedures and the two-sample log-rank test. In this figure, we consider the case $\kappa=1$ so that we can also meaningfully compare tests in which Weibull MLEs are used with those in which exponential MLEs are used. Both suffice the standard assumption discussed in Section 3 and hence are consistent and asymptotically normal. However, as the shape parameter has to be estimated for a Weibull distribution, the exponential estimate might be more efficient and thus yield a higher power of the associated test. However, our results reveal that these expectations are not being met. The corresponding lines in Figure 4 overlap almost completely. This is not only the case for the presented results with $n_{B}=100$ and $\pi=0.25$, but also for the other scenarios which are presented in the Supplementary Material.
Compared to the two-sample log-rank test, there are no clear advantages resulting from the parametric assumption. If Wu’s variance estimation is used, there are advantages of a few percentage points for small allocation ratios. However, for allocation ratios close to 1, this effect is reversed. These observations are consistent with the conservativeness of the two-sample log-rank test for small allocation ratios observed in the previous section and described in [11], and with the conservativeness of the corrected one-sample log-rank test for allocation ratios close to 1, which was also observed above.
Results for other settings support these findings an can also be found in the Supplementary Material.
@@ 7 Discussion
As we have demonstrated here, a considerable inflation of the type I error rate occurs if the one-sample log-rank test is to be interpreted as a test of the practically highly relevant null hypothesis $H_{0}$. This is the case as the calculation of the test statistic does not take the estimation uncertainty of the reference curve into account. With the analytical derivation and a simulation study, we were able to demonstrate that we can decompose the variance of the test statistic under the hypothesis $H_{0}$ into two parts, one of which corresponds to the well-known variance estimator for the one-sample log-rank test and the other addresses the uncertainty in the reference curve. As can be seen from formula (7), the allocation ratio between the historical control cohort and the experimental cohort is of fundamental importance. While the one-sample log-rank test in its current form can still play an important role as a helpful tool, particularly in the early stages of treatment development, especially in rare diseases or in paediatrics, we would like to draw attention to the fact that there are ways of taking a more critical statistical view of the role of the reference curve. For example, the test statistics developed here could be considered as an additional sensitivity analysis to the one-sample log-rank test.
Of course, it can be argued that the resulting comparison is ultimately a two-group comparison between the historical cohort and the study cohort, which could also be achieved with a two-sample log-rank test. By making an additional parametric assumption about the distribution of the endpoint in the historical control group, it seems conceivable at first glance that the resulting tests are more efficient and allow for gains in terms of power in comparison with the two-sample log-rank test, which does not require such assumptions. However, our simulation study revealed that no systematic gains can be achieved in a proportional hazards scenario. In comparison with the two-sample log-rank test there is a slight advantage in cases where the historical control cohort is relatively large in comparison with the experimental group. For groups of similar size, on the other hand, there is a slight disadvantage. These findings are consistent with the observed conservativeness of the respective methods described in Section 6.2.1. Analogously to the uncorrected one-sample procedures, the choice of the estimator for the variance component $V_{1}$ plays a crucial role for the corrected tests in terms of power. We refer to [5] for a more extensive discussion of this issue.
In summary, based on our simulation results, it can be said that conducting the two-sample log-rank test is preferable if the necessary data is available and a test of the hypothesis $H_{0}$ is desired instead of a comparison with a fixed reference curve. However, it should be noted that our simulations are limited to proportional hazards scenarios, in which the two-sample log-rank test is known to be very efficient. If the one-sample log-rank test shall still be applied for the primary analysis, our proposed method could be still be used as an additional sensitivity analysis or as a tool to determine how much of the total variance of the test statistics can be attributed to the uncertainty inherent in the estimate of the reference curve.
@@
Going beyond the standard one-sample log-rank test, the techniques presented and evaluated here are in general also applicable to variations of the one-sample log-rank test. This includes weighted tests [4], and resulting combination tests [17] as well as multi-stage procedures [13, 16].
@@ Appendix A Derivation of the asymptotic distribution
$-n_{B}^{1/2}\left(\mathbb{E}[\Lambda(\hat{\boldsymbol{\theta}}_{A},t\wedge X_{B})|\hat{\boldsymbol{\theta}}_{A}]-\mathbb{E}[\Lambda(\boldsymbol{\theta}_{A},t\wedge X_{B})]\right)$ (17)
@@ 50 Refer to caption Refer to caption
B.2 Type I error rate (Exponential and Weibull MLEs)
$n_{B}$ two-sided rates one-sided rates
25 Refer to caption Refer to caption
50 Refer to caption Refer to caption
100 Refer to caption Refer to caption
200 Refer to caption Refer to caption
Supplementary Figure S4: Two- and one-sided empirical rejection rates of the null hypothesis $H_{0}$ with exponentially distributed data for procedures based on exponential and Weibull MLEs for four different sample sizes in the experimental cohort in dependence of the allocation ratio.
B.3 Power
$n_{B}$ power
25 Refer to caption
50 Refer to caption
100 Refer to caption
200 Refer to caption
Supplementary Figure S5: Power to reject the null hypothesis $H_{0}$ for Weibull-distributed data with shape parameter $\kappa=0.5$ and hazard ratio $\omega=0.8$ for four different sample sizes in the experimental cohort in dependence of the allocation ratio.
$n_{B}$ power (excl. procedures with exponential MLEs) power (incl. procedures with exponential MLEs)
25 Refer to caption Refer to caption
50 Refer to caption Refer to caption
100 Refer to caption Refer to caption
200 Refer to caption Refer to caption
Supplementary Figure S6: Power to reject the null hypothesis $H_{0}$ for Weibull-distributed data with shape parameter $\kappa=1$ and hazard ratio $\omega=0.8$ for four different sample sizes in the experimental cohort in dependence of the allocation ratio. The right column also includes procedures based on exponential MLEs
$n_{B}$ power
25 Refer to caption
50 Refer to caption
100 Refer to caption
200 Refer to caption
Supplementary Figure S7: Power to reject the null hypothesis $H_{0}$ for Weibull-distributed data with shape parameter $\kappa=2$ and hazard ratio $\omega=0.8$ for four different sample sizes in the experimental cohort in dependence of the allocation ratio.